# W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction

### 29 Nov 2023 v1

@@ 3.3 Orientation Correction
Table 1: Comparison with SOTA methods on distorted datasets in camera coordinate. Some of the evaluation results on HuMMan [3] and SPEC-MTP [19] are sourced from  [43], while the evaluation results on AGORA [35] are obtained from its official website https://agora-evaluation.is.tuebingen.mpg.de/. Including the following tables, $*{}^{*}$ represents backbone is ResNet50 [11], ${}^{\dagger}$ means VitPose-B [47, 48], and ”P” denotes the pure parameter module. Note we do not perform any fine-tuning on SPEC-MTP. The baseline structure is kept the same as [52], but its training dataset and evaluation settings are the same as ours.
3.4 Training Paradigm
[…]
$\mathcal{L}_{f}=MSE\left(f,\tilde{f}\right)$ (7)
where $\tilde{f}=\sqrt{W^{2}+H^{2}}$ serves as pseudo focal length labels to ensure focal length predictor outputs coarse but relatively reasonable results. Note it is only applied in stage 1.
[…]
In the first two stages, training is conducted entirely in camera coordinate. In the third stage, we need to equip the model to reconstruct the human body in world coordinate. The pre-trained regressors, focal length predictor, feature extractor, and CamCalib are frozen at this stage. We only need to train the OrientCorrect composed of three FC layers. We exclusively use the 3D joint, mesh vertices, and SMPL parameters in world coordinate for supervision.
4 Experiment
4.1 Implement Details
We train three models based on [52] for evaluation. Two of them utilize a hybrid regression module, using 431 sparse vertices for downsampling. The distinction lies in using different backbones, including ResNet50 [11] and VitPose-B [47]. Additionally, we train an extra model using a pure parameter regression module, employing 67 mocap markers for downsampling, with the backbone being VitPose-B. For training, we utilize datasets such as Human 3.6m [12], 3DPW [42], MPII3D [31], MPII [1], AGORA [35], COCO [28], SPEC-SYN [19], and HuMMan [3]. Some of the datasets are augmented with pseudo-labels generated by CamCalib, EFT [13], or CLIFF [25]. Please refer to the supplementary materials for detailed structural variations and training settings of our different models. Our model can simultaneously provide results in camera and world coordinate. But if results in world coordinate are not needed or camera rotation is known, the CamCalib and OrientCorrect can be discarded. The focal length predictor can also be discarded if mesh projection is not required.
@@ 4.2 Evaluation Results
We conduct evaluations on multiple benchmarks. Our model’s performance presentation is divided into camera and world coordinates in the main text. Our model is also evaluated on conventional benchmarks such as 3DPW [42] and h36m [12], yielding outstanding results too. Supplementary materials are available for further reference. Below are the main results and analysis.
@@
Table 2: Comparison with SOTA methods on SPEC-MTP [19]. The first six models in the table are all methods that output results in camera coordinate, so there are two results. The ”a” in ”a/b” indicates the result calculated directly using the output in the camera coordinate. The ”b” is the result calculated after the output has been calibrated by camera rotation predicted by CamCalib. The latter two models can directly get the results in world coordinate without calibration, so there is only one result.
Figure 5: Qualitative results on SPEC-MTP [19]. ”Input” refers to input images. ”Camera” represents the output in the camera coordinate. ”World” represents the output in world coordinate.
Camera Coordinate In our study, we choose three challenging datasets containing distorted images to showcase the superiority of weak-supervised camera calibration under complex conditions. As depicted in Tab. 1, our method surpasses the performance of most existing methods on the challenging synthetic dataset AGORA [35]. Each of our models demonstrates unique strengths. The pure regression model achieves higher accuracy in joint localization, and the hybrid regression module excels in capturing mesh details. Compared to baseline [52], our improvement is quite substantial. As depicted in Fig. 4, the W-HMR outperforms SPEC [19] in handling occlusions, rare body rotations, and body intersections. Although the joint accuracy does not exhibit significant improvement on the indoor dataset HuMMan [3], we achieve an 11% improvement in PVE, even surpassing the model-free method FastMETRO [4]. The most remarkable performance is observed on the real distorted dataset SPEC-MTP. We significantly elevate the SOTA results on SPEC-MTP to new heights.
@@
World Coordinate We utilize SPEC-MTP [19] as the primary evaluation dataset. Due to its only test set, specific fine-tuning is not feasible. It adopts a shooting method different from traditional datasets, resulting in unique rotation and severe distortion. Therefore, it poses a significant challenge and is particularly suitable for demonstrating OrientCorrect’s strength. To ensure fairness, we only evaluate the W-HMR with a pure parameter regression module. W-HMR easily outperforms existing methods, including SPEC. It is important to note that W-HMR does not rely solely on higher reconstruction accuracy. Upon comparing our results calibrated by CamCalib with those corrected by OrientCorrect, it is evident that our module is effectively functioning. As illustrated in Fig. 5, the reconstructed human body pose appears implausible in world space when the camera coordinate is not aligned with the world coordinate. However, our OrientCorrect effectively corrects the orientation of the human body and maintains its reasonableness in world coordinate. Our model can accurately perceive and correct both slight and significant camera rotation. Quantitative and qualitative results both demonstrate that our method is simpler and more efficient than SPEC, which will be further substantiated in subsequent ablation study.
4.3 Ablation Study
Our ablation study is primarily divided into two parts. The first part focuses on ablating our model’s camera calibration to demonstrate our approach’s effectiveness in enhancing reconstruction accuracy. Specifically, the main idea is to verify the validity of the global information for the camera calibration. The second part involves ablating different global orientation recovery methods. We will show the limitations of previous methods and demonstrate the robustness of W-HMR. Please refer to supplementary materials for more ablation study.
AGORA
[…]
$\surd$ 82.4 88.6 75.8 81.5
[…]
$\surd$ $\surd$ 80.7 87.7 73.4 79.8
$\surd$ $\surd$ $\surd$ 79.8 87.6 72.6 79.7
[…]
Camera Calibration We select PyMAF[52] as our baseline and consistently use ResNet50 as the backbone throughout our ablation study, with 431 sparse vertices as the downsampling points. The training and evaluation settings are all kept the same. From the Tab 3, it is evident that the most significant improvements stem from the addition of the vertex regressor and global information. The inclusion of the vertex regressor enhances the model’s ability to reconstruct body mesh details accurately. But, global information not only contributes independently but also aids in the functioning of camera calibration. A comparison between ”V” and ”V+F” reveals that solely adding the focal length predictor only has a noticeable impact on vertex accuracy. However, comparing ”V+G” and ”V+F+G” demonstrates that the presence of global information leads to more substantial improvements when combined with the focal length predictor. This finding further supports our idea that global information is crucial not only for recovering body orientation but also for predicting focal lengths and correcting distortions. In summary, with the camera calibration assistance, W-HMR achieves a decent reconstruction accuracy, which provides a solid foundation for human body recovery in world space.
Model SPEC-SYN
[…]
GraphCMR [21] 181.7/181.5 86.6 219.8/218.3
SPIN [20] 165.8/161.4 79.5 194.1/188.0
PartialHumans [38] 169.3/174.1 88.2 207.6/210.4
I2L-MeshNet [33] 169.8/163.3 82.0 203.2/195.9
HMR [14] 128.7/96.4 55.9 144.2/111.8
PyMAF [52] 126.8/88.0 48.7 136.7/118.0
${}^{\dagger}$ 127.8/106.2 46.2 137.8/132.6
SPEC [19] 74.9(124.3) 54.5(71.8) 90.5(147.1)
${}^{\dagger}$(spec-like) 68.2(185.0) 43.6(63.3) 76.4(208.7)
${}^{\dagger}$ 82.1(118.7) 46.1(66.6) 93.3(133.9)
Table 4: The ablation study on body orientation correction. ”spec-like” refers to using a similar method like SPEC to predict global pose directly. The numbers in brackets are the evaluation results on SPEC-MTP. This is the reason why we don’t choose the spec-like method. Its impressive results are based on overfitting and sacrificing generalization.
Orientation Correction We select two methods for evaluation. The first method is similar to SPEC. SPEC concatenates features and the camera parameters obtained from CamCalib, then inputs them into a regressor to directly recover the human body in world coordinate. As demonstrated in Tab. 4, the spec-like method significantly outperforms previous methods. However, we attribute this performance to overfitting. The model relies on massive parameters to memorize the global body orientation in the training set of SPEC-SYN, subsequently achieving excellent results on the test set. This is supported by the inferior performance of spec-like methods, such as results on SPEC-MTP. This indicates the susceptibility of such methods to overfitting, potentially leading to the misconception that the model has captured environmental information. What’s worse is that this design destroys the model’s original learned knowledge, making the overall reconstruction less precise. In contrast, our method offers a simpler structure, requires less training data, and exhibits greater robustness. Furthermore, other models based on camera coordinate can easily adopt our approach to achieve global pose recovery.
5 Conclusion
We proposed the first model capable of reconstructing human bodies accurately and reasonably in both camera and world coordinates. To achieve this, we designed a weak-supervised camera calibration method. Our model can generate appropriate focal lengths based on body distortion information, eliminating the dependence on background information and focal length labels. Additionally, we introduced a new orientation correction module that is simpler and more robust compared to previous methods. With a carefully designed training paradigm and model framework, we achieved SOTA performance across multiple datasets in both coordinates. We hope that our work will contribute to the study of human mesh recovery in world space.
Supplementary Material
6 About Datasets
In our experiments, we utilize the following datasets: Human 3.6M [12], 3DPW [42], MPII [1], COCO [28], MPI-INF-3D [31], AVA  [9], AIC [45], Insta [15], AGORA [35], SPEC-SYN [19], SPEC-MTP [19], and HuMMan [3]. To evaluate our model, we conduct experiments on the distorted datasets AGORA, HuMMan, and SPEC-MTP. Furthermore, following traditional practices[44, 52], we also evaluate our model on 3DPW, Human 3.6M, and MPII-INF-3D. For the ablation experiments, we employ AGORA’s evaluation platform as a standard due to its diverse scenes, characters, and shooting methods. We believe that the evaluation results obtained on AGORA are more representative. For human recovery in world space, We additionally perform ablation study on SPEC-SYN to validate the robustness of our method. Following are details about each dataset.
3DPW’ [42] full name is ”3D Human in the Wild.” It is an outdoor dataset that captures actor information using motion capture devices. This dataset is known for its difficulty and has been widely used by researchers to demonstrate the robustness of their methods. Following past conventions, we train two models, one without 3DPW for training and another using 3DPW for training, to evaluate our model’s performance.
Human 3.6M [12] is widely used as the standard dataset to assess the estimation of 3D human pose and the recovery of human mesh. To reduce data redundancy, the original videos are down-sampled from 50 to 10 fps, resulting in 312,188 frames available for training and 26859 frames for evaluation. Following established protocols, we utilize five subjects (S1, S5, S6, S7, S8) for training and two subjects (S9, S11) for evaluation.
AVA [9], AI Challenger[45] and InstaVariety [15] are three huge wild datasets that lack real 3D labels. However, in the Human4D [8], researchers utilized VitPose [47] to obtain 2D keypoints and VitDet [24] to generate bounding boxes for human targets. Based on these 2D pseudo labels, they employed ProHMR [22] to acquire pseudo SMPL labels. Subsequently, we further optimize these pseudo-labeled datasets by selecting higher-quality samples for training our model.
[…]
$H_{g}\times W_{g}$ 21$\times$21 8$\times$8 24$\times$18 9$\times$7
[…]
Table 5: This table illustrates the changes in the size of input images and features when using different backbones (ResNet50 or VitPose-B) and employing different downsampling methods (sparse vertices or markers). As for the symbols, please refer to Sec 8.1.
7 Other Losses
Since we employ a hybrid regression module, we obtain 431 sparse vertices, 1723 coarse vertices, and 6840 complete vertices from the vertex regressor. Therefore, we design a triple vertex loss function, denoted as
where $\beta$ is shape parameters in SMPL and $\theta$ is pose parameters which is represented as 9D rotation matrix.
To enhance the model’s ability to focus on human regions and extract refined body information, we include two additional sub-tasks for auxiliary supervision inspired by [52].
First, let’s delve into the IUV map, which is integral to the IUV loss. The IUV map, as defined in [10], uses ”I” to indicate the body part to which a pixel belongs and ”UV” to represent the 2D coordinates on the surface of the body part. This mapping establishes dense correspondences from 3D mesh to 2D plane. We employ three 2D convolution layers to predict the IUV map based on the feature map $\varPhi_{3}$ and render the ground truth online. The loss function is defined as
where $\odot$ means masking, which indicates only the foreground pixels are applied. The IUV loss focuses solely on supervising the 2D information, neglecting the depth information, which is essential for capturing distorted details. To address this limitation, we introduce an additional depth map loss. We generate ground truth depth maps by rendering online. A 2D convolution is used to predict the depth map based on the feature map $\varPhi_{3}$. The loss is computed as follows:
where we use the reciprocal of the depth for normalization.
8 Implement Details
8.1 Pixel-aligned Feature Extractor
We follow the feature extraction method of [52] but use different sampling points and backbones. When a cropped image $I\in\mathbb{R}^{H_{c}\times W_{c}}$ is inputted into the encoder, it generates an initial feature map $\varPhi_{0}\in\mathbb{R}^{C\times H_{0}\times W_{0}}$. Subsequently, this feature map $\varPhi_{0}$ undergoes further processing through a series of deconvolution layers to yield high-resolution features $\left\{\varPhi_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}\right\}_{i=1}^{3}$. For the first high-resolution feature map $\varPhi_{1}$, a $H_{g}\times W_{g}$ grid is defined to sample finer point-wise features. However, for the second and third feature maps, we project vertices of the mesh obtained from the previous step onto the feature map to sample point-wise features. In summary, we extract spatial features based on the following formula.
[…]
where $P_{i}$ represents the projected 2D point for sampling, $bs(\cdot)$ denotes bilinear sampling, and $fc(\cdot)$ refers to FC layers for reducing channel size. Finally, $\oplus$ signifies the concatenation of individual point-wise features $\left\{\phi_{i,n}^{p,fc}\in\mathbb{R}^{C_{i}^{fc}}\right\}_{n=1}^{N_{i}}$ into a one-dimensional feature $\phi_{i}^{s}\in\mathbb{R}^{C_{i}^{s}}$, where $C_{i}^{s}=N_{i}\times C_{i}^{fc}$. $fc$ indicates the point-wise features has been through $fc(\cdot)$ and $N_{i}$ is the number of points for downsampling.
8.2 Hybrid Regression Module
For the spatial features $\{\phi_{i}^{s}\}_{i=1}^{3}$, we have corresponding regression models. We utilize parameter regressors for the first two spatial features to obtain the initial meshes. These regressors take spatial features $\phi_{i}^{s}$, global information $glob\_info$, and former SMPL parameters $\varTheta_{i-1}$ as inputs and output SMPL parameters $\varTheta_{i}$.
Figure 6: $\phi_{3}^{s}$ is first inputted into an FC layer to reduce its dimension to match the dimensions of other tokens. The 3D coordinates of the vertices are concatenated with their corresponding point-wise feature, forming a vertex query.
The parameter regressors consist of several FC layers and iteratively produce the final result like
Note $\varTheta_{0}$ is set as the mean $\overline{\varTheta}$ following [14]. As for the final regressor, we have two versions in total. The first version still involves the parameter regressor. The second version utilizes a vertex regressor to obtain a finer human body mesh. After obtaining a coarse human mesh using the first two regressors, we introduce a model-free vertex regressor based on the transformer. As depicted in Fig. 6, this regressor takes spatial features $\phi_{3}^{s}$ and point-wise features $\phi_{3}^{p}$ as input. It’s important to note that no ”$fc$” exists in this case, indicating that these features do not undergo channel reduction. To reduce network complexity and adapt to the SMPL model, instead of directly recovering the 6890 vertices defined by SMPL, we recover 431 sparse vertices and gradually restore them to 6890 vertices using two linear layers.
8.3 Summary
We experiment with different backbones and downsampling points to validate our method and explore the optimal model structure. As a result, there are corresponding variations in the sizes of input images and features. We summarize these variations and create Tab 5 to provide an overview of these changes. The training process is conducted with a batch size of 128 and a fixed learning rate of 0.00005 on three RTX 3090 GPUs until convergence, taking approximately one week. Following [52, 18], we apply standard data augmentation techniques such as rotation, flipping, scaling, translation, noise, and occlusion. As for hyperparameters, $\lambda_{1}=\lambda_{2}=\lambda_{3}=300$, $\lambda_{4}=\lambda_{5}=\lambda_{6}=15$, $\lambda_{7}=0.06$,$\lambda_{8}=60$, $\lambda_{9}=0.03$ and $\lambda_{10}=\lambda_{11}=\lambda_{12}=0.05$.
9 About Metrics
We evaluate the reconstruction accuracy using 3D mesh vertices and joints. Specifically, for AGORA, we employ four metrics: MPJPE, MVE, NMJE, and NMVE. MPJPE stands for Mean Per Joint Position Error, which measures the accuracy of joint positions. For AGORA, it calculates the error for the first 24 SMPL joints after aligning with the root joint. Following [52, 20], we utilize 17 joints for MPI-INF-3D, while the other datasets employ a consistent set of 14 keypoints. MVE, or Mean Vertex Error, measures the error of all SMPL mesh vertices after aligning the root vertices. For other datasets except AGORA, PVE (Per Vertex Error) is computed without alignment. These metrics help assessing the accuracy of the reconstructed human bodies. NMJE (Normalized Mean Joint Error) and NMVE (Normalized Mean Vertex Error) are obtained by normalizing MPJPE and MVE using the F1 score. They provide an overall evaluation of the method’s performance, particularly in handling challenging scenarios. For specific calculation methods, please refer to [35].
As for the metrics in world coordinate, we follow SPEC [19] using W-MPJPE, PA-MPJPE, and W-PVE. PA-MPJPE (Procrustes-aligned Mean Per Joint Position Error) exists as a metric that purely evaluates the accuracy of body reconstruction. It removes the effect of unknown camera rotation by Procrustes-aligned and has therefore been adopted by many previous works. The calculation of W-MPJPE and W-PVE is not changed. The only difference is that the labels used are adjusted to the ground truth in world coordinate. We, including SPEC, believe that these metrics are more representative of the model’s capabilities in real-world applications.
10 More Experimental Results
@@
Table 6: Comparison with SOTA Methods on 3DPW [42]. The column ”w/ 3DPW” indicates whether the model has been trained using the 3DPW dataset.
@@
Traditional Benchmarks Following previous methods [52, 44], we conduct evaluation on 3DPW [42], H36M [12], and MPII-INF-3D [31]. Specifically, we evaluate the performance of models trained with and without the 3DPW train set on the 3DPW test set. As shown in Tab 6, we achieve remarkable improvements on this challenging in-the-wild dataset, surpassing previous works easily. Even compared to the SOTA method HybrIK [23], W-HMR still exhibits significant advancements. However, in Tab 7, W-HMR’s performance is relatively less impressive in the case of controllable datasets, although it remains competitive with these SOTA methods. Controllable datasets refer to datasets obtained under laboratory conditions, typically characterized by limited variations in actors and scenes, which are prone to overfitting. HuMMan is also an example of such a dataset. So, it becomes difficult for our advantages to be fully demonstrated.

### 25 Mar 2024 v3

@@ 3.2 About Datasets
AVA [9], AI Challenger[46] and InstaVariety [15] are three huge wild datasets that lack real 3D labels. However, in the Human4D [8], researchers utilized VitPose [48] to obtain 2D keypoints and VitDet [24] to generate bounding boxes for human targets. Based on these 2D pseudo labels, they employed ProHMR [22] to acquire pseudo SMPL labels. Subsequently, we further optimize these pseudo-labeled datasets by selecting high-quality samples for training our model.
3.3 About Metrics
We evaluate the reconstruction accuracy using 3D mesh vertices and joints. Specifically, for AGORA, we employ four metrics: MPJPE, MVE, NMJE, and NMVE. MPJPE stands for Mean Per Joint Position Error, which measures the accuracy of joint positions. For AGORA, it calculates the error for the first 24 SMPL joints after aligning with the root joint. Following [54, 20], we utilize 17 joints for MPI-INF-3D, while the other datasets employ a consistent set of 14 keypoints. MVE, or Mean Vertex Error, measures the error of all SMPL mesh vertices after aligning the root vertices. PVE (Per Vertex Error) is computed without alignment for other datasets except AGORA. These metrics help assess the accuracy of the reconstructed human bodies. NMJE (Normalized Mean Joint Error) and NMVE (Normalized Mean Vertex Error) are obtained by normalizing MPJPE and MVE using the F1 score. They provide an overall evaluation of the method’s performance, particularly in handling challenging scenarios. For specific calculation methods, please refer to [35].
As for the metrics in world coordinate, we follow SPEC [19] using W-MPJPE, PA-MPJPE, and W-PVE. PA-MPJPE (Procrustes-aligned Mean Per Joint Position Error) exists as a metric that purely evaluates the accuracy of body reconstruction. It removes the effect of unknown camera rotation by Procrustes-aligned and has therefore been adopted by many previous works. The calculation of W-MPJPE and W-PVE is the same as MPJPE and PVE. The only difference is that the labels used are adjusted to the ground truth in world coordinate. We, including SPEC, believe these metrics represent the model’s capabilities in real-world applications.
@@ 3.4 Evaluation Results
We conduct massive experiments on multiple benchmarks in a variety of formats. The evaluation results can be categorized into three main parts. Firstly, we do experiments on the distorted dataset (AGORA, HuMMan, and SPEC-MTP) in camera and world coordinates, respectively, to demonstrate the effectiveness of W-HMR on distorted images. Secondly, we also do rich experiments on traditional benchmarks (3DPW, H36M, and MPI-INF-3D) to compare the past SOTA methods to demonstrate our model’s superiority. Finally, we provide rich qualitative results to visualize the performance of W-HMR.
Camera Coordinate In our study, we choose three challenging datasets containing distorted images to showcase the superiority of weak-supervised camera calibration under complex conditions. As depicted in Table 2, our method surpasses the performance of most existing methods on the challenging synthetic dataset AGORA [35]. Each of our models demonstrates unique strengths. The pure regression model achieves higher accuracy in joint localization, and the hybrid regression module excels in capturing mesh details. Compared to baseline [54], our improvement is quite substantial. Note the baseline’s backbone, training datasets and evaluation settings are the same as $*{}^{*}$. As depicted in Fig. 5, our W-HMR outperforms SPEC [19] in handling occlusions, rare body rotations, and body intersections. Although the joint accuracy does not exhibit significant improvement on the indoor dataset HuMMan [3], we achieve the closest performance compared to the latest method Zolly in PVE, even surpassing the model-free method FastMETRO [4]. The most remarkable performance is observed on the real distorted dataset SPEC-MTP. We significantly elevate the SOTA results on SPEC-MTP to new heights.
Table 3: Comparison with SOTA methods on SPEC-MTP [19]. The first six models in the table are all methods that output results in camera coordinate, so there are two results. The "a" in "a/b" indicates the result calculated directly using the output in the camera coordinate. The "b" is the result calculated after the output has been calibrated by camera rotation predicted by CamCalib. The latter two models can directly get the results in world coordinate without calibration, so there is only one result.
@@
World Coordinate We utilize SPEC-MTP [19] as the primary evaluation dataset. Due to its only test set, specific fine-tuning is not feasible. It adopts a shooting method different from traditional datasets, resulting in unique rotation and severe distortion. Therefore, it poses a significant challenge and is particularly suitable for demonstrating OrientCorrect’s strength. To ensure fairness, we only evaluate the W-HMR with a pure parameter regression module. W-HMR easily outperforms existing methods, including SPEC. It is worth noting that the superior performance of W-HMR in world coordinates is not entirely dependent on higher reconstruction accuracy. Perhaps you noticed that there are two Ours-Ps in Table 3. This is because W-HMR can generate results in two coordinates simultaneously. By comparing these two sets of results, we can conclude that OrientCorrect is significantly superior to the traditional method of directly using predicted camera rotation for correction. Upon comparing our results calibrated by CamCalib with those corrected by OrientCorrect, it is evident that our module is effectively functioning. As illustrated in W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction, the reconstructed human body pose appears implausible in world space when the camera coordinate is not aligned with the world coordinate. However, our OrientCorrect corrects the orientation of the human body and maintains its reasonableness in world coordinate. Our model can accurately perceive and correct both slight and significant camera rotation. Quantitative and qualitative results both demonstrate that our method is simpler and more efficient than SPEC, which will be further substantiated in subsequent ablation studies.
Table 4: Comparison with SOTA Methods on 3DPW [42]. The column "w/ 3DPW" indicates whether the model has been trained using the 3DPW dataset. We bold both the best results with and without 3DPW.
@@
Traditional Benchmarks Following previous methods [54, 45], we conduct evaluation on 3DPW [42], H36M [12], and MPII-INF-3D [31]. Specifically, we evaluate the performance of models trained with and without the 3DPW train set on the 3DPW test set. As shown in Table 4, we achieve remarkable improvements on this challenging in-the-wild dataset, surpassing previous works easily. Even compared to the SOTA method Zolly, W-HMR still exhibits significant advancements when trained without the 3DPW train set, which shows W-HMR’s robustness. In Table 5, W-HMR’s performance is relatively less impressive in the case of controllable datasets, although it remains competitive with these SOTA methods. Controllable datasets refer to datasets obtained under laboratory conditions, typically characterized by limited variations in actors and scenes, which are prone to overfitting. HuMMan in Table 2 is also an example of such datasets. So, it becomes difficult for our advantages to be fully demonstrated.
Qualitative Results More visualizations are generated to give a more comprehensive reference. Also, we welcome running our codes on your samples. Back to qualitative results, from Fig. 7, we can see that our model is able to recover the human body very well when facing various scenes with various characters in multiple poses. We purposely show the results from side angles and in world space. In the past, many methods could also make the reconstructed mesh match the original image perfectly, but it would not look very good if we changed the visual angle. As shown in Fig. 7, W-HRM doesn’t suffer from this issue and performs well in every condition.
Figure 6: Qualitative results on HuMMan [3]. "Input" refers to input images of the model. "Overlay" means we project the recovered mesh back onto the original images. "Cam" and "World" show the output in camera and world coordinates, respectively.
Methods based on camera coordinate often fail to meet the needs of applications in world space. It may not be noticeable from Fig. 7 because they fit the weak-perspective camera model and the assumption of $R_{c}=I$. But from Fig. 6, the drawbacks of the traditional method are undeniable. The reconstructed human body is entirely lopsided in world coordinate. Instead, our OrientCorrect module can correct the pose well to achieve a reasonable pose recovery in world coordinate. This method has a wide range of applications, allowing the model to reconstruct a proper human body when facing in-the-wild images. Before W-HMR, there were also methods like SPEC that explored human mesh recovery in world space. As a pioneering work, SPEC opened up a new way of thinking for subsequent research, but the performance was not robust enough. In Fig. 8, we can see that SPEC is not as accurate as W-HMR in any coordinate, showing that our method cleverly avoids the error accumulation of SPEC and achieves all-around transcendence.
Figure 8: Qualitative comparison with SPEC on SPEC-MTP [19]. Here are some examples: W-HMR shows better mesh-image alignment and higher orientation accuracy in world space.
3.5 Ablation Study
Our ablation study is primarily divided into three parts. The first part focuses on ablating our model’s camera calibration to demonstrate our approach’s effectiveness in enhancing reconstruction accuracy. More specifically, we explore the interplay between the focal length predictor and other designs. The second part involves ablating different global orientation recovery methods. We show the limitations of SPEC and demonstrate the robustness of OrientCorrect. In the third part, we do more ablation studies about some functional designs but not our core contribution.
[…]
AGORA
[…]
87.3 92.6 78.6 83.3
$\surd$ 82.4 88.6 75.8 81.5
$\surd$ $\surd$ 81.2 89.3 74.7 82.2
$\surd$ $\surd$ 80.7 87.7 73.4 79.8
$\surd$ $\surd$ $\surd$ 79.8 87.6 72.6 79.7
Camera Calibration We select PyMAF[54] as our baseline and consistently use ResNet50 as the backbone throughout our ablation study, with 431 sparse vertices as the downsampling points. The training and evaluation settings are all kept the same. From the Table 6, it is evident that the most significant improvements stem from adding the vertex regressor and global information. Including the vertex regressor enhances the model’s ability to reconstruct body mesh details. However, global information not only contributes independently but also aids in the functioning of camera calibration. A comparison between "V" and "V+F" reveals that solely adding the focal length predictor only has a noticeable impact on vertex accuracy. However, comparing "V+G" and "V+F+G" demonstrates that the presence of global information leads to more substantial improvements when combined with the focal length predictor. This finding further supports our idea that global information is crucial for recovering body orientation, predicting focal lengths, and correcting distortions. In summary, with the camera calibration assistance, W-HMR achieves a decent reconstruction accuracy, which provides a solid foundation for human body recovery in world space.
Table 7: The ablation study on body orientation correction. "spec-like" refers to using a similar method like SPEC to predict global pose directly. The numbers in brackets are the evaluation results on SPEC-MTP. This is the reason why we don’t choose the spec-like method. Its impressive results are based on overfitting and sacrificing generalization.
Model SPEC-SYN
[…]
GraphCMR [21] 181.7/181.5 86.6 219.8/218.3
SPIN [20] 165.8/161.4 79.5 194.1/188.0
PartialHumans [38] 169.3/174.1 88.2 207.6/210.4
I2L-MeshNet [34] 169.8/163.3 82.0 203.2/195.9
HMR [14] 128.7/96.4 55.9 144.2/111.8
PyMAF [54] 126.8/88.0 48.7 136.7/118.0
${}^{\dagger}$ 127.8/106.2 46.2 137.8/132.6
SPEC [19] 74.9(124.3) 54.5(71.8) 90.5(147.1)
${}^{\dagger}$(spec-like) 68.2(185.0) 43.6(63.3) 76.4(208.7)
${}^{\dagger}$ 82.1(118.7) 46.1(66.6) 93.3(133.9)
Orientation Correction We select two methods for evaluation. The first method is similar to SPEC. SPEC concatenates features and the camera parameters obtained from CamCalib, then inputs them into a regressor to directly recover the human body in world coordinate. As demonstrated in Table 7, the spec-like method significantly outperforms previous methods. However, we attribute this performance to overfitting. The model relies on massive parameters to memorize the global body orientation in the training set of SPEC-SYN, subsequently achieving excellent results on the test set. This is supported by the inferior performance of spec-like methods on SPEC-MTP. This indicates such method is prone to overfitting, which can lead to the misconception that the model can capture environmental information. What’s worse is that this design destroys the model’s original learned knowledge, making the overall reconstruction less precise. In contrast, our method offers a simpler structure, requires less training data, and exhibits greater robustness. Furthermore, other models based on camera coordinate can easily adopt our approach to achieve global pose recovery.
Table 8: Ablation study results on SPEC-MTP. The weights are from SPIN and are not fine-tuned. ${}^{\dagger}$ indicates using orientation output by OrientCorrect of pre-trained W-HMR. $*{}^{*}$ indicates HMR with an adaptive OrientCorrect module.
[…]
HMR 173.6 75.4 206.6
${}^{\dagger}$ 149.6 75.4 179.4
$*{}^{*}$ 133.9 75.4 157.9
In order to verify the portability of OrientCorrect, we do related experiments, and the results are displayed in Table 8. To make the experiments more convincing, we choose HMR, a widely adopted model, as the baseline. We design two porting methods. One is to directly use the pre-trained OrientCorrect from W-HMR to predict the body rotation in the world space and then combine it with the pose and shape output from HMR to generate the final mesh. It can be seen that it brings nearly 14% improvement. From Table 3, it can be seen that using the camera rotations output from SPEC’s CamCalib to correct brings only about 10% boost. Moreover, the most impressive is the second porting method. We utilize the features and body orientation output by HMR to replace the original input part of OrientCorrect. In other words, we train an OrientCorrect adapted to HMR. The effect is pronounced, and the reconstruction accuracy is again improved by one grade. To avoid misunderstandings, you may think OrientCorrect is not as good as CamCalib or SPEC because the results in Table 8 are not as good as those of HMR and SPEC in Table 3. This is since we are based on different baselines. We use the pre-trained weights from SPIN directly and do not do any tuning, so the final results should be worse.
Design Details To find the optimal design for our model, we conduct further ablation experiments, as shown in Table 9. Firstly, we test different downsampling methods. Evidently, using markers for downsampling leads to better learning of human body features. Here, "markers" refer to mocap markers. The improved performance might be because mocap markers are a better proxy for inferring human body information [56]. However, since using mocap markers conflicts with the vertex regressor, we only employ markers in the pure parameter regression.